The Earth's magnetic field is 10amp/m.How much magnetic energy is stored in 1m^3 of the atmosphere because of this field? Estimate also the electric energy density associated with this magnetic field assuming plane waves to be present.(pH=pE=6.28×10^-5J/s.m^3)
1
Expert's answer
2021-07-15T15:21:30-0400
"p_H=0.5(4\pi \cdot10^{-7})(10)^2(1)=\frac{W}{m^3}"